OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form.

OpenStudy (anonymous):

y - 5 = 3(x - 2)

OpenStudy (anonymous):

So you know that slope is equal to 3 (m) Plug it into the equation (y -y1) = m (x - x1) as (2,5) = (x1, y1) and m = 3

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

y-5=3(x-2) y-5=3x-6 3x-y-6+5= 3x-y-1= this is the equation of point slope form